# Stability and Controllability of Revenue Systems via the Bode Approach

### v1

@@ 1 Introduction
A pacing system creates a budget spending plan (referred to the desired spend velocity) based on the advertising product being managed. This plan typically spans a single day or multiple days over the course of an advertising campaign. Equipped with one or more feedback loops, the pacing system monitors actual spending performance (the actual spend velocity) and actively adjusts it in near real-time to align with the spending plan. By minimizing the error between the desired and actual spend velocities, the system ensures targets are achieved [1, 2].
@@ 4.1 Laplace Transformation
$ZOH\cdot G(s)\cdot H(s)=\frac{1-e^{T_{ps}}}{s}\cdot W_{n}\cdot\frac{1}{1+sT_{f}}$ (6)
@@
$ZOH\cdot G(s)\cdot H(s)=\frac{W_{n}T_{ps}}{1+sT_{f}}\sum\limits_{n=1}^{9}\frac{\left({-sT_{ps}}\right)^{n-1}}{n!}$ (8)
@@ 6.2 Simulation of the Compensated Plant
PI compensator: In the compensated pacing system, the open-loop transfer function is given by $G_{c}(s)\cdot ZOH\cdot G(s)\cdot H(s)$. It is easy to prove that P control alone is insufficient to stabilize the system (proof in the Appendix), and D control is undesirable as it amplifies feedback noises and thereby reducing the signal-to-noise (S/N) ratio. In the Practitioner method, we select a PI compensator and perform a grid search for parameters that satisfy the Bode stability criteria. To prevent integrator windup, the integrator is limited to the range $[0,0.5]$ and is only activated when the lumped output is within the range $(0,1)$. Nine combinations of $K_{p}$ and $K_{i}$ with their stability performance can be found in Table 1. We can see that a stronger integrator (larger $K_{i}$) must match with a higher gain (larger $K_{p}$) to maintain stability. However, $K_{p}$ will find its limitation to stabilize the system once $K_{i}$ goes beyond a certain point. Their corresponding plots are shown in Figure 5.
@@ A.2 Exponential Smoother
Appendix B Proof of Instability using P Control Only
This section demonstrates the instability of using P control alone. To illustrate this, we calculate the P-compensated GM by substituting $s$ with $j\omega$:
$\parallel G_{c}(s)\cdot ZOH\cdot G(s)\cdot H(s)\parallel=K_{p}W_{n}\frac{\parallel 1-e^{-j\omega T_{ps}}\parallel}{\parallel 1+j\omega T_{f}\parallel}=1.0$ (24)
Note that to mathematically prove instability, we must use the original formulation of the ZOH rather than the Taylor approximation. We then convert the exponential expression into trigonometric functions using Euler’s formula $e^{ix}=\cos\left(x\right)+i\cdot\sin\left(x\right)$.
$\begin{array}[]{l}{K_{p}W_{n}\frac{\sqrt{(1-cos(-\omega T_{ps}))^{2}+(sin(-\omega T_{ps}))^{2}}}{\sqrt{1+\omega^{2}{T_{f}}^{2}}}}\\ {=K_{p}W_{n}\frac{\sqrt{1-2cos\left(\omega T_{ps}\right)}}{\sqrt{1+\omega^{2}{T_{f}}^{2}}}=1.0}\end{array}$ (25)
For $0<\omega<\frac{1}{T_{ps}}$, since $01-2\cos(\omega T_{ps})<$ 0, there is no real-valued $K_{p}$ that can compensate the plant to achieve a positive gain margin.
Appendix C Tested Ad Sets Info

### v2

@@ 1 Introduction
A pacing system creates a budget spending plan (referred to as the desired spend velocity) based on the advertising product being managed. This plan typically spans a single day or multiple days over the course of an advertising campaign. Equipped with one or more feedback loops, the pacing system monitors actual spending performance (the actual spend velocity) and actively adjusts it in near real-time to align with the spending plan. By minimizing the error between the desired and actual spend velocities, the system ensures targets are achieved [1, 2].
@@ 4.1 Laplace Transformation
$ZOH\cdot G(s)\cdot H(s)=\frac{1-e^{-sT_{ps}}}{s}\cdot W_{n}\cdot\frac{1}{1+sT_{f}}$ (6)
@@
$ZOH\cdot G(s)\cdot H(s)=\frac{W_{n}T_{ps}}{1+sT_{f}}\sum\limits_{n=2}^{10}\frac{\left({-sT_{ps}}\right)^{n-1}}{n!}$ (8)
@@ 6.2 Simulation of the Compensated Plant
PI compensator: In the compensated pacing system, the open-loop transfer function is given by $G_{c}(s)\cdot ZOH\cdot G(s)\cdot H(s)$. Since P control alone is insufficient to stabilize the system due to missing high-frequency noise rejection, and D control is undesirable as it amplifies feedback noises and thereby reducing the signal-to-noise (S/N) ratio. In the Practitioner method, we select a PI compensator and perform a grid search for parameters that satisfy the Bode stability criteria. To prevent integrator windup, the integrator is limited to the range $[0,0.5]$ and is only activated when the lumped output is within the range $(0,1)$. Nine combinations of $K_{p}$ and $K_{i}$ with their stability performance can be found in Table 1. We can see that a stronger integrator (larger $K_{i}$) must match with a higher gain (larger $K_{p}$) to maintain stability. However, $K_{p}$ will find its limitation to stabilize the system once $K_{i}$ goes beyond a certain point. Their corresponding plots are shown in Figure 5.
@@ A.2 Exponential Smoother
Appendix B Tested Ad Sets Info